# See, Imagine, Plan: Discovering and Hallucinating Tasks from a Single Image

### v2

@@ 3.1.3 Task Proposal
We query VLM to propose meaningful, diverse tasks, each with a one-sentence task description. Instead of directly querying VLM for task proposal, we employ a hybrid approach that integrates role-play and object-based initialization. In the role-play scenario, we prompt VLM to envision itself as a robotic/human hand working in the scene to perform household tasks. For the object-based initialization, we guide VLM to sequentially focus on each identified object within the scene. When the scene contains more than one identified objects, VLM is instructed to suggest two tasks emphasizing interactions between the manipulating object and any of the detected objects, and an additional task focused solely on the manipulating object. If only one object is detected, VLM is directed to propose a task involving just that object. This strategy guarantees a broad spectrum of task suggestions, ensuring comprehensive object engagement.
@@
To further tailor the task proposals, we impose specific constraints, directing VLM to consider the practical affordances of objects while encouraging creative assumptions (e.g., a bowl’s capacity to hold water) and potential interactions (e.g., transferring water from a cup into a bowl). Additionally, we delineate clear boundaries by excluding tasks that entail the assembly or disassembly of objects, functionality tests, or the involvement of imaginary objects, thereby focusing on feasible and meaningful tasks.
@@
3.2 Reconstructing and Understanding 3D Image Scene
Figure 3: Our method of partial 3D scene reconstruction. (a) full representation of foreground objects and background as a plane, (b) the reconstructed scene from the active camera view, (c) the original input image.
Our method for partial reconstruction of the 3D scene focuses on full representation of foreground objects while simplifying the inpainted background as a plane. The reconstructed scene is captured by a pinhole camera in the 3D environment, as shown in Fig. 3(a). Given that our framework is designed to accommodate input images captured by users using arbitrary cameras, we make the assumption that the camera intrinsic matrix is unknown. Our partially reconstructed scene demonstrates high visual alignment with the input image, see Fig. 3(b) and (c).
3.2.1 Single-View 3D Object Reconstruction and Depth Estimation
The first step of reconstructing the 3D image scene involves reconstructing 3D object models and estimating depth. We feed occlusion-free object masks, $\{M_{k}^{of}\}_{k=1}^{K}$, to an off-the-shelf single-view 3D reconstruction model to acquire object 3D models, $\{O_{k}\}_{k=1}^{K}$. We also feed the original map $I$ to an off-the-shelf depth estimation model to acquire depth image, denoted as $I_{dep}$.
3.2.2 Camera 6D Pose Estimation and Object Scale Initialization
Camera 6D Pose Estimation   Single-view 3D reconstruction model reconstructs mesh $O_{i}$ of object $i$ at the 3D origin, in the coordinate frame set by the input mask $M_{i}^{of}$, and captures $O_{i}$ by a pinhole camera. This camera, with 6D pose $P_{c}=[R_{c}\mid t_{c}^{origin}]$, captures $O_{i}$’s canonical pose within the image. Thus, we can restore all objects’ canonical poses across all images by identifying $P_{c}$.
We use One-2-3-45++ [38], which provides the camera pose. For models without this information, we develop an efficient method to determine the pose by comparing object masks with rendered 3D model templates, inspired from matching-based 6D pose estimation works [34, 46, 9, 41]. Details in Appendix. G.
For pinhole camera to focus on the background plane center within the reconstructed 3D scene, we backproject the 2D center of the inpainted background $I_{bg}$ using a camera intrinsic matrix $C_{\text{Mat}}$, and adjust the translation vector $t_{c}^{origin}$ in the camera pose $P_{c}$. $C_{\text{Mat}}$ is arbitrary, with the only condition of principal point coordinates, $c_{x}$ and $c_{y}$, set to image $I$ center. Focal lengths $f_{x}$ and $f_{y}$ are set to $\0\epsilon\in\mathbb{R}\backslash\{0\}$ because backprojecting the image center results in $u=c_{x}$ and $v=c_{y}$. Thus, our camera intrinsic matrix and adjusted, final camera pose are:
[…]
Setting Background Plane   The 6D pose of the background plane $O_{bg}$ (visually identical to $I_{bg}$), denoted as $P_{bg}=[R_{bg}\mid t_{bg}]$, matches the camera’s rotation $R_{c}$ as its orientation aligns with the camera. The position $t_{bg}$ is shifted along the camera’s directional vector ($v_{c}=-R_{c}\times[0,0,1]$), ensuring fitting precisely within the camera’s view based on the camera’s focal length and sensor width.
Object Scale Initialization   3D reconstruction models typically standardize objects within a normalized sphere [71, 45, 13, 39, 38]. To set the initial scale for $O_{i}$, we multiply the contour length of its occlusion-free 2D mask $M_{i}^{of}$ by a predefined scale factor $\delta_{init}$, tailored to model’s normalized space dimensions:
$S_{i}^{init}=\text{arcLength}(\text{findContour}(M_{i}^{of}))\times\delta_{init}.$ (3)
However, this method overlooks objects’ depth, inaccurately assuming uniform distance from the camera (e.g., moving an apple from close to the camera to a distant corner reduces its contour length). To correct this, we refine the object scales after determining their 6D poses, detailed in the subsequent Sec. 3.2.3.
3.2.3 Object 6D Pose Estimation and Object Scale Calibration
Object 6D Pose Estimation   Our goal is to position object 3D models into the reconstructed 3D scene without visual discrepancies and ensure accurate depth. Because a significant portion of depth estimation model [69, 51, 4, 5] is trained on depth datasets with depth data determined by sensors [21, 55] and stereo matching [70, 15, 63, 64, 67], we assume that the predicted normalized depth is the perpendicular distance to the camera plane, instead of a straight line from the object to the camera lens [30]. We start by determining the 2D centers of each occlusion-free object mask, $\{M_{k}^{of}\}_{k=1}^{K}$, using center of mass. Then, we calculate the corresponding object 3D centers on the background plane $O_{bg}$, denoted as $\{t_{k}^{bg}\}_{k=1}^{K}$. This process involves initially mapping the 2D centers onto $O_{bg}$ upon its importation into the scene and then transforming these points by multiplying the background plane’s pose $P_{bg}$. Similarly, denote the camera plane as $O_{c}$ with pose $P_{c}=[R_{c}\mid t_{c}]$ and each object 3D center on $O_{c}$ as $\{t_{k}^{c}\}_{k=1}^{K}$. For an object $i\in[1,K]$, we perform raycasting from $t_{i}^{c}$ on the camera plane to $t_{i}^{bg}$ on the background plane, and its distance $d_{i}$ to move from $t_{i}^{c}$ towards $t_{i}^{bg}$ is written as:
$d_{i}=\left|k_{init}\times I_{dep}(\text{center}(M_{i}^{of})\right|.$ (4)
Here, $k_{init}$ represents the depth scaling factor that adjusts the normalized depth values to the appropriate scale. Because the pinhole camera is used to capture the 3D scene, we then raycast from the camera lens $t_{c}$ to $t_{i}^{bg}$ with distance calculated based on $d_{i}$ such that the perpendicular distance moved from the camera plane towards $t_{i}^{bg}$ remains unchanged. The 3D coordinate of object $i$ is:
$t_{i}=t_{c}+\frac{d_{i}}{\left|t_{i}^{bg}-t_{i}^{c}\right|}\times(t_{i}^{bg}-t_{c}).$ (5)
The rotation matrix $R_{i}$ of the 6D pose of object $i$, $P_{i}=[R_{i}\mid t_{i}]$, is set to identity matrix because the reconstructed 3D model $O_{i}$ is aligned in its canonical pose.
Object Scale Calibration   After integrating all 3D object models $\{O_{k}\}_{k=1}^{K}$ into the 3D scene, each with a pose $\{P_{k}=[R_{k}\mid t_{k}]\}_{k=1}^{K}$ and an initial scale $\{S_{k}^{init}\}_{k=1}^{K}$, we refine their scales to accurately reflect depth variations. Through the lens of the pinhole camera with pose $P_{c}$, we render binary masks for each object and evaluate the length of their contour lines relative to their occlusion-free masks. The adjusted, final scale of object $i$ can be expressed as:
where $M_{i}^{rend}$ is the rendered mask of object $i$.
Finding Object Principle Axes   We determine the principal axes (x-axis, y-axis, and z-axis) of each object, which is essential for our novel 3D motion planning method, explained in Sec. 3.3.1. Using the minimal oriented bounding box (OBB), we encapsulate each object in the smallest possible box, aligning the object’s principal axes with the box’s dimensions. OBB employs principal component analysis on the object’s point cloud to identify directions of maximum, second greatest, and least variance, which naturally correspond to the object’s orientation along its length, width, and height. This approach is straightforward and effective, especially since our input images primarily consist of rigid objects.
3.2.4 Image and Spatial Context Understanding
With accurate camera pose, background plane pose, and object poses and scales, we can reconstruct the 3D scene, denoted as $V_{0}$, from the input image $I$. Before using VLM for motion planning within the 3D scene based on proposed tasks, we conduct in-context learning to familiarize VLM with the reconstructed 3D environment. We feed $V_{0}$ with visualized object axes (see Fig. 2B). Visualizing 3D spatial information is pivotal in improving VLMs’ understanding of 3D spatial contexts derived from 2D images, validated by 3DAxiesPrompts [35]. To augment VLM’s understanding, we also feed a paragraph describing the objects’ poses, sizes, and principal axes in physical units, alongside their spatial relationships. This information is derived entirely from our reconstructed 3D scene. As a concrete example for the left image on Fig. 2:
‘‘Obj 0 spatial context:
- 3D center: [-21.0, 101.0, 4.0] cm
- Local x-axis (towards ‘right’): [0.8637, 0.293, -0.41]
- Local y-axis (towards ‘back_left’): [-0.5943, 0.7879, 0.1614]
- Local z-axis (towards ‘front’): [-0.1606, -0.7822, 0.6019]
Obj 0 size: 29.55 cm x 26.91 cm x 13.19 cm (WxDxH)
Obj 0 - Closest per direction: right: Obj 1 (29.16 cm); down: Obj 4 (15.72 cm) …’’
3.3 Planning and Task Execution within Reconstructed 3D Scene
3.3.1 Axes-Constrained Motion Planning through Waypoints
@@
We introduce a novel method to guide VLM to conduct motion planning within a 3D scene based on a proposed task by planning motion waypoints along the manipulating object’s principal axes. More specifically, we define four types of manipulations that VLM can use:
@@
3.3.2 Trajectory Generation and Optimization
Handling VLM Planning Discrepancies   The waypoints generated by VLM are typically accurate and practical. Nonetheless, there are instances where the waypoints suggested by VLM lead to collisions as determined by the RRT* planner. This discrepancy is less about VLM’s misunderstanding of the objects’ sizes and their spatial relationship and more about the precision level of the waypoints, which may not match the exacting standards of the RRT* planner’s outcomes. To resolve this, we implement Gaussian sampling around the initially planned waypoints whenever a collision is detected. The sampling strategy is guided by a predefined set of geometric rules. Please see our Appendix. G for details.
Trajectory Optimization   Finally, to ensure our trajectory is natural and smooth, we linearly interpolate rotation and interpolate translation using cubic spline. The final trajectory for proposed task $q\in[1,Q]$ is denoted as $V^{q}=\{V_{r}^{q}\}_{r=0}^{R}$. Each pose in the final trajectory $V^{q}$ is rendered as a frame in the rendering engine, where $V_{0}^{q}=V_{0}$, the rendering of the reconstructed 3D scene.
4 Experiments
Figure 4: Dataset statistics. Our dataset presents 51 scenes—13 from NOCS and 38 captured from varied perspectives—featuring a wide range of object categories, quantities, and a diverse set of tasks, task videos, and planned trajectories.
Implementations   We use ChatGPT-4 [47] as our VLM. The language-guided segmentation model is Language Segment-Anything [43] and the repainting model is LaMa [60]. We use One-2-3-45++ [38] for single-view 3D reconstruction and Depth Anything [69] for depth estimation. For partial 3D scene reconstruction, we use Blender [16] as the 3D software. The pinhole camera focal length and sensor width are set to 80 mm and 36 mm, with $\delta_{init}=2.5e-4,k_{init}=1$. All inference is run on 1 NVIDIA A10 GPU with 24GB RAM.
Dataset   As we are the first to propose zero-shot task hallucination, there lacks a pre-existing dataset for evaluation. Therefore, we craft a diverse evaluation dataset by combining self-captured photos and scenes from the NOCS dataset [62]. We capture 38 photos using an iPhone 12 Pro Max. From NOCS, we use its real-world part of both the training and test sets, which encompasses 13 distinct scenes. Since each scene is a video sequence, we randomly extract one frame as our data. We standardize all image dimensions by resizing all to $640\times 480$.
Dataset Statistics   Our dataset covers diverse scenes (e.g., office, kitchen, bathroom), and features a rich diversity of object categories (116) and quantities (185), with each image containing $1-7$ objects and $1-3$ tasks proposed for each object (278 tasks/task videos/planned trajectories in total). The dataset’s diversity is further enhanced by the variety of perspectives from which the images are captured or selected (e.g., frontal, top-down, side views). This deliberate choice of diverse angles, both in our own image capturing process and through the random extraction of frames from NOCS, aims to simulate a realistic and challenging array of scenes for evaluation. See Fig. 4 for statistics and visuals.
4.1 Task Diversity
@@
We evaluate the diversity of the proposed tasks in terms of semantic meaning using Self-BLEU and the embedding similarity [74] following RoboGen [66], where lower scores mean higher diversity. We also compare with previous reinforcement learning (RL) benchmarks. From Table. 1, ours method generates most diverse tasks as our pipeline is open to all scenes with no constraint.
4.2 Qualitative Demonstration
Figure 5: Qualitative results. With diverse input scenes and proposed tasks, our model produces high-quality task execution visualizations with geometric awareness that aligns with the task descriptions. Zoom in for better view.
We present qualitative examples spanning various scenes and tasks, illustrating our pipeline’s capability to produce diverse, visually appealing, and accurate hallucinated task videos for a wide range of input images, as shown in Fig. 5. Additional examples are available at our project page.
Figure 6: Comparison with 3DIT [18] and Runway Gen-2 [52] on rotation and translation tasks. 3DIT fails to maintain consistency and precision across interpolated frames. Runway produces unpredictable artifacts. Our task videos are geometric-aware, photo-realistic with fine-grained and consistent object motion. Zoom in for better view.
We also compare with our baselines, SOTA 3D-aware image editing and video generation models, 3DIT [18] and Runway Gen-2 [52], as shown in Fig. 6. We limit comparisons to rotation and translation tasks to avoid bias against 3DIT and Runway in more nuanced context-dependent tasks (e.g., slicing an apple), where their performances fall short. For 3DIT, we generate and connect frames with our planned trajectories. For Runway, we prompt it with task descriptions.
4.3 Human Preference Evaluation: User Study
Table 2: User study. We ask 25 users to rate (scale $1-5$) and rank (scale $1-3$) for 5 rotation and 5 translation tasks, and rate (scale $1-5$) for 10 complex context-dependent tasks. All results take average. Note that our user study size is similar to those representative works such as ControlNet [73] and Prompt-to-Prompt [23].
Rotation Translation Context-Dependent Rating $\uparrow$ Ranking $\downarrow$ Rating $\uparrow$ Ranking $\downarrow$ Rating $\uparrow$ 3DIT [18] 1.93 2.34 1.90 2.27 - Runway [52] 1.55 2.62 1.43 2.68 - Ours 4.58 1.04 4.43 1.05 4.29
We rely on human preference evaluation as one of our quantitative metrics. We let users rate and rank the rotation and translation task videos generated by 3DIT [18] and Runway Gen-2 [52] against our own in terms of both video quality and task description alignment. For these complex context-dependent tasks, we instead ask users to judge the videos’ authenticity and execution relative to human action, and to encapsulate their perception of the action in our videos with a single sentence. The usage of sentence description will be explained in the subsequent Sec. 4.4. Table. 2 reveals the details and results.
4.4 Understanding Hallucination
Machine Understanding
Table 3: Results for machine understanding (classification) on all 278 task videos.
Machine
Raw Acc $\uparrow$ 0.974
Fal-Pos Rate $\downarrow$ 0.363
True Acc $\uparrow$ 0.611
We assess the interpretability of our generated task videos from a machine’s perspective using SOTA video understanding model, Video-LLaVA-7B [33]. We use two approaches: binary classification and descriptive generation. For classification, we feed the model with the task descriptions generated by VLM and ask question (is the video $\ldots$?). In generation, we prompt Video-LLaVA-7B to articulate its interpretation of the tasks depicted in our videos. To quantify the correspondence between the model’s perception and the tasks, we use OpenCLIP cosine similarity score [14].
However, we find that even SOTA video understanding model shows limited performance. To assess false positive rate in classification, we deliberately misalign the sequence of generated task videos with their corresponding task descriptions, expecting a theoretical accuracy of 0%. Contrary to expectations, Video-LLaVA-7B reports a false positive rate of 36.3%. To adjust for this anomaly, we subtract this rate from the model’s raw accuracy for correctly aligned video-task pairs. This method, while unconventional, provides a more fair and reasonable evaluation of machine video understanding, underscoring the current challenges faced by video understanding models in accurately interpreting complex video content. Results in Table. 3.
Table 4: Results for machine understanding (generation) on all 278 task videos and human understanding, where 25 users write descriptions for 5 videos.
Machine Human
OpenCLIP $\uparrow$ 0.636 0.823
Human Understanding   We aim to understand human perception of our generated task videos by asking participants to provide a one-sentence description. We then evaluate the alignment between these descriptions and the ground truth task descriptions proposed by VLM, using OpenCLIP [14]. Table. 4 reveals a high degree of alignment. Intriguingly, it appears humans understand our hallucinated videos more accurately than machines do. We hypothesize this discrepancy stems from the limitations of current video understanding models, whereas humans draw on their prior experiences for a deeper comprehension.
4.5 Task Proposal Ablation Study
Table 5: Task proposal ablation study. We use all 278 proposed tasks.
Ours Ours w/o obj-init Ours w/o role-play Ours w/o both
Involved Objects Number $\uparrow$ 185 144 185 146
Self-BLEU $\downarrow$ 0.273 0.280 0.276 0.282
Embedding Similarity $\downarrow$ 0.143 0.156 0.141 0.159
We aim to understand if our approach of task proposal, employing role-play and object-based initialization, yields the most diverse and feasible tasks by modifying the prompt to exclude these techniques. In addition to using Self-BLEU and the embedding similarity, we also count the number of objects involved across all 278 proposed tasks.
The results in Table. 5 indicate that our method yields the most diverse tasks. Evaluating on all 278 proposed tasks shows a slight rise in Self-BLEU and a decrease in embedding similarity compared to the 106 tasks sampled (Table. 1). This difference arises because more task descriptions lead to repetitive structures and wording, yet the semantics grow more varied. Excluding object-based initialization reduces diversity by causing the VLM to overlook certain objects. Semantic meaning of tasks cannot reflect feasibility. Instead, we show instances where VLM generates impractical tasks without role-play in Appendix. H. We find that the impact of removing role-play is more complex, affecting task diversity less (Table. 5) but reducing task feasibility.
5 Discussion
We proposed a framework for zero-shot task hallucination. From the observation of a single image, our framework is able to discover diverse tasks and hallucinate task executions as geometric-aware and photo-realistic videos with fine-grained and consistent object motion. We see our work as a step towards equipping machines with the profound human capacity, in which humans can recall past knowledge and anticipate future scenarios through visual perception. Our generated task videos and planned trajectories could spark exciting applications, such as letting robots automatically explore and perform tasks in new environments.
Limitations   Not all generated task videos achieve the highest quality. We identified key factors that contribute to the failure cases. Please see Appendix. I.
Appendix
F Project Page
All code, data, and additional videos are available at our project page: https://dannymcy.github.io/zeroshot_task_hallucination/.
G Additional Details of Method
G.1 Matching-Based Camera 6D Pose Estimation
For single-view 3D reconstruction models which do not provide the 6D pose, $P_{c}=[R_{c}\mid t_{c}^{origin}]$, of the pinhole camera, we develop an efficient method to determine the pose by comparing object masks with rendered 3D model templates, inspired from matching-based 6D pose estimation works [34, 46, 9, 41].
$m_{i}^{A,h}&=\text{Hu}(\text{findContour}(\text{crop}(M_{i}^{of}))),\;\;\;\;\;\;pa_{i}^{A}=\text{sum}(\text{resize}(\text{crop}(M_{i}^{of})))\\ m_{i,j}^{B,h}&=\text{Hu}(\text{findContour}(\text{crop}(T_{j}^{i}))),\;\;\;\;\;\;\;\;pa_{i,j}^{B}=\text{sum}(\text{resize}(\text{crop}(T_{j}^{i})))\\ \mathcal{L}(M_{i}^{of},T_{j}^{i})&=\alpha\left|1-\frac{\min(pa_{i}^{A},pa_{i,j}^{B})}{\max(pa_{i}^{A},pa_{i,j}^{B})}\right|+\beta\sum_{h=1}^{7}\left|\frac{1}{\text{sgn}(m_{i}^{A,h})\cdot\log(m_{i}^{A,h})}-\frac{1}{\text{sgn}(m_{i,j}^{B,h})\cdot\log(m_{i,j}^{B,h})}\right|\end{aligned}$
This dual approach allows for a comprehensive comparison that incorporates both the geometric configuration and the scale of the object representations. The best-matched template can be found by $\arg\min_{j=1}^{J}\mathcal{L}(M_{i}^{of},T_{j}^{i})$, and we set its corresponding camera pose to be the 6D pose of our pinhole camera, denoted as $P_{c}=[R_{c}\mid t_{c}^{origin}]$.
The camera now captures the object’s canonical view, centered at the 3D origin. Subsequently, we adjust the camera’s translation component to focus on the center of the inpainted background $I_{bg}$, resulting in $P_{c}=[R_{c}\mid t_{c}]$, as detailed in our main paper.
G.2 Sampling Rules for VLM Planning Discrepancies
Figure 7: Feasibility comparison of the proposed tasks with and without role-play.

### v5

@@ C.2 Discovering and Planning for Robotics Tasks from a Single Image
Task Proposal   We query VLM to propose meaningful, diverse tasks, each with a one-sentence task description. Instead of directly querying VLM for task proposal, we employ a hybrid approach that integrates role-play and object-based initialization. In the role-play scenario, we prompt VLM to envision itself as a robotic/human hand working in the scene to perform household tasks. For the object-based initialization, we guide VLM to sequentially focus on each identified object within the scene. When the scene contains more than one identified objects, VLM is instructed to suggest two tasks emphasizing interactions between the manipulating object and any of the detected objects, and an additional task focused solely on the manipulating object. If only one object is detected, VLM is directed to propose a task involving just that object. This strategy guarantees a broad spectrum of task suggestions, ensuring comprehensive object engagement.
@@
To further tailor the task proposals, we impose specific constraints, directing VLM to consider the practical affordances of objects while encouraging creative assumptions (e.g., a bowl’s capacity to hold water) and potential interactions (e.g., transferring water from a cup into a bowl). Additionally, we delineate clear boundaries by excluding tasks that entail the disassembly of objects, functionality tests, or the involvement of imaginary objects, thereby focusing on feasible and meaningful tasks.
@@
Axes-Constrained Motion Planning through Waypoint   We introduce a novel method to guide VLM to conduct motion planning within a 3D scene based on a proposed task by planning motion waypoints along the manipulating object’s principal axes. More specifically, we define four types of manipulations that VLM can use:
@@
In practice, the quality of motion planning by VLMs can be enhanced using various prompting techniques. One such technique is chain-of-thought (CoT) [63], where another LLM guides the VLM to plan each axes-constrained sparse waypoint step by step.
Dataset Statistics   See Fig. 8 for statistics and visuals. We standardize all image dimensions by resizing all to $640\times 480$.
Figure 8: Dataset statistics. Our dataset presents 51 scenes—13 from NOCS and 38 captured from varied perspectives—featuring a wide range of object categories, quantities, and a diverse set of tasks and planned trajectories.
More Qualitative Demonstration   We present more qualitative examples spanning various scenes and tasks, as shown in Fig. 9.
Figure 9: More qualitative examples. With diverse input scenes and proposed tasks, our framework produces 3D trajectories with geometric awareness that aligns with the task descriptions. Zoom in for better view.
Experiment on Human Understanding
Table 9: Results for machine understanding (generation) on 278 task executions and human understanding, where 25 users write descriptions for 10 tasks.
Machine Human
OpenCLIP $\uparrow$ 0.636 0.823
We aim to understand human perception of our generated task executions by asking participants to provide a one-sentence description. We then evaluate the alignment between these descriptions and the ground truth task descriptions proposed by VLM, using OpenCLIP [15, 43]. Table. 9 reveals a high degree of alignment. Intriguingly, it appears humans understand our task executions more accurately than machines do (Table. 7). We hypothesize this discrepancy stems from the limitations of current video understanding models, whereas humans draw on their prior experiences for a deeper comprehension.
Experiment on Task Diversity   We evaluate the diversity of the proposed tasks in terms of semantic meaning using Self-BLEU and the embedding similarity [75, 48, 49] following RoboGen [61], where lower scores mean higher diversity. We also compare with previous reinforcement learning (RL) benchmarks. From Table. 10, ours method generates most diverse tasks as it is open to all scenes with no constraint.
@@
C.3 Trajectory Generation Implementation Details